Derive the equation (23) at page 42 of Unit 2, which represents the polar equation of a conic when the directrix L corresponding to a focus F is taken to the right of F.
1
Expert's answer
2020-05-19T13:48:40-0400

Let F be a fixed point, the focus, and let L be a fixed line, the directrix, in a plane. A conic section, or conic, is the set of all points P in the plane such that 


"\frac{PF}{PL}=e,\ e>0"

where 

e is a fixed positive number, called the eccentricity.

If e=1,

the conic is a parabola.

If e<1, 

the conic is an ellipse.

If e>1, 

the conic is a hyperbola.

By locating a focus at the pole, all conics can be represented by similar equations in the polar coordinate system. In each of these equations,

(r,θ) is a point on the graph of the conic.

e is the eccentricity.



p is the distance between the focus (located at the pole) and the directrix.

For a conic with a focus at the origin, if the directrix is "x=\pm p,"

where 

p is a positive real number, and the eccentricity is a positive real number e,

the conic has a polar equation


"r=\frac{
e⋅p}{1±cosθ}"

For a conic with a focus at the origin, if the directrix is "y=\pm p,"

where p is a positive real number, and the eccentricity is a positive real number e,the conic has a polar equation


"r=\frac{
e⋅p}{1±sinθ}"

Given that the directrix L

L corresponding to a focus F

F is taken to the right of F.

Directrix "L: x=p"




The distance from the focus to the point P in polar is just r.

PF=r

The distance from the point P to the directrix is  x=p is "PL=p-r\cos{\theta}" .

 Then

"e=\frac{PF}{PL}=\frac{r}{p-r\cos{\theta}}"

Solve for r


"e.p-e .\cos\theta.r=r"

"r=\frac{
e⋅p}{1+cosθ}"

is the polar equation.